# Asymptotic stability equals exponential stability—while you twist your eyes

### v1

@@ Appendix A Continuation and the Conley index
The existence of a homotopy on $M^{n}$ is not particularly insightful by itself. More interesting is to consider classification up to homotopy, e.g., Hopf’s celebrated degree theorem states (in its simplest form) that all continuous maps from $\mathbb{S}^{m}$ to itself, are completely classified up to homotopy, via their (topological) degree, e.g., see [ref:milnor65, p. 51].
The question that we try to address is motivated by better understanding to what extent Hopf’s theorem extends to qualitatively equivalent classes of dynamical systems.
The most fruitful framework in this regard is Conley’s theory [ref:conley1978isolated]. It is outside the scope of this note to cover these ideas, but we briefly highlight it to elucidate the central question of this work.
In the context of semiflows, Conley’s index theory can be set up as follows [ref:rybakowski1987homotopy, Ch. 1]. Given a semiflow $\varphi$ on $M^{n}$, then, $S\subseteq M^{n}$ is said to be an isolated (positively) invariant set when there is a compact set $K\subseteq M^{n}$, called an isolating neighbourhood, such that
$S=\mathrm{I}(K,\varphi):=\{p\in K:\varphi^{t}(p)\in K\,\forall t\geq 0\}\subseteq\mathrm{int}(K).$
Now, a pair of compact sets $(N,L)\subset M^{n}\times M^{n}$ is said to be an index pair for $S$ when
  1. $({I}.i)$
    $S=\mathrm{I}(\mathrm{cl}(N\setminus L),\varphi)$ and $N\setminus L$ is a neighbourhood of $S$;
  2. $({I}.ii)$
    $L$ is positively invariant in $N$; and
  3. $({I}.iii)$
    $L$ is an exit set for $N$.
Then, the (homotopy) Conley index of $S$ is the homotopy type of the pointed (quotient) space $(N/L,[L])$. Importantly, the Conley index is independent of the choice of index pair. Towards a computational theory, one usually works with the homological definition instead.
Now, if some $N$ can be chosen to be an isolating neighbourhood throughout a homotopy $[0,1]\ni s\mapsto\varphi(\cdot;s)$, then, the Conley index is preserved along that homotopy and we speak of a continuation (of the Conley index). The existence of the homotopy implies continuation, but to what extent does an equivalent Conley index imply the existence of an index preserving homotopy? This is precisely the starting point of this note and this is why we speak of a “converse question”.

### v2

@@ Appendix A Continuation and the Conley index
We aim to contribute to better understanding to what extent Hopf’s degree theorem [ref:milnor65, p. 51] extends to equivalence classes of dynamical systems. The most fruitful framework in this regard is Conley’s theory [ref:conley1978isolated] and we briefly highlight his index theory to elucidate our central question.
In the context of semiflows, Conley’s index theory can be set up as follows [ref:rybakowski1987homotopy, Ch. 1]. Given a semiflow $\varphi$ on $M^{n}$, then, $S\subseteq M^{n}$ is said to be an isolated (positively) invariant set when there is a compact set $K\subseteq M^{n}$, called an isolating neighbourhood, such that $S=\mathrm{I}(K,\varphi):=\{p\in K:\varphi^{t}(p)\in K\,\forall t\geq 0\}\subseteq\mathrm{int}(K)$. Now, a pair of compact sets $(N,L)\subset M^{n}\times M^{n}$ is said to be an index pair for $S$ when (i) $S=\mathrm{I}(\mathrm{cl}(N\setminus L),\varphi)$ and $N\setminus L$ is a neighbourhood of $S$; (ii) $L$ is positively invariant in $N$; and (iii) $L$ is an exit set for $N$. Then, the Conley index of $S$ is the homotopy type of the pointed space $(N/L,[L])$. This index is independent of the choice of index pair. Now, if some $N$ can be chosen to be an isolating neighbourhood throughout a homotopy $[0,1]\ni s\mapsto\varphi(\cdot;s)$, then, the Conley index is preserved along that homotopy and we speak of continuation (of the index). The existence of the homotopy implies continuation, but to what extent does an equivalent Conley index imply the existence of an index preserving homotopy? This is the starting point of this note and this is why we speak of a “converse question”.
Appendix B Added in proof
Recently, Kvalheim and Sontag provided a generalized Hartman-Grobman theorem that allows for resolving open problem 2 [ref:kvalheim2025global, Thm. 2] (e.g., first homotope to a smooth flow), see also [ref:jongeneel2025HG]. Moreover, Kvalheim addressed the extension to vector fields with outstanding generality [ref:kvalheim2025differential], see [ref:kvalheim2025differential, §8.4] for an example that affirms our hopes at the end of Section 2.